### v1

# Low-Dimension-to-High-Dimension Generalization And Its Implications for Length Generalization

Abstract
Low-Dimension-to-High-Dimension (LDHD) generalization is a special case of Out-of-Distribution (OOD) generalization, where the training data are restricted to a low-dimensional subspace of the high-dimensional testing space. Assuming that each instance is generated from a latent variable and the dimension of the latent variable reflects the problem scale, the inherent scaling challenge in length generalization can be captured by the LDHD generalization in the latent space. We theoretically demonstrate that LDHD generalization is generally unattainable without exploiting prior knowledge to provide appropriate inductive bias. Specifically, we explore LDHD generalization in Boolean functions. We verify that different architectures trained with (S)GD converge to min-degree interpolators w.r.t. different independent sets. LDHD generalization is achievable if and only if the target function coincides with this inductive bias. Applying the insights from LDHD generalization to length generalization, we explain the effectiveness of CoT as changing the structure latent space to enable better LDHD generalization. We also propose a principle for position embedding design to handle both the inherent LDHD generalization and the nuisances such as the data format. Following the principle, we propose a novel position embedding called RPE-Square that remedies the RPE for dealing with the data format nuisance.
1 Introduction
The field of learning to reason has gained significant popularity in the machine learning community due to its impressive performance in reasoning tasks such as natural language processing [16, 17], mathematics [8, 11], coding [28], symbolic logic [1, 9], and planning [30, 21]. Reasoning problems inherently require generalization beyond the training distribution, a challenge known as Out-of-Distribution (OOD) generalization [13, 27]. One of the most significant challenges in OOD generalization is scaling, also known as length generalization [3, 29], where models trained on small-scale instances must generalize to large-scale instances. Length generalization is crucial because the size of sample spaces often increases exponentially with the complexity of reasoning problems, leading to intractable sample complexity and computational costs for models that do not achieve length generalization.
Numerous works have investigated the length generalization problem in various tasks and proposed various techniques that help scaling in practice, including modifications to model architectures [20, 11, 12], transformations in data formats [14, 31], prompt engineering for Large Language Models (LLMs) Wei et al. [22], Feng et al. [7]. While some techniques work uniformly well across a wide class of problems, many are fragile and even ad-hoc, applicable only to specific problems with certain formats [32]. This can be largely attributed to the mismatch between the inherent problem scale and the formal scale in the input domain (e.g., the length of the input string in language models). We further elaborate the statement in Example 1.
Example 1.
Consider the addition learning problem in the language domain, where the input is a string. For an $N$-digit-plus-$N$-digit (written as $N$-addition) addition of two numbers $x$ and $y$, where $x=x_{N-1}\dots x_{0}$, $y=y_{N-1}\dots y_{0}$, $x_{i},y_{i}\in\{0,\dots,9\}$ ($x_{N-1}>0$ or $y_{N-1}>0$), consider two formats of the instance: the Aligned and Reverse Format (ARF), where the instance is represented as “$\mathtt{x_{0}\dots x_{n-1}+y_{0}\dots y_{n-1}=}$”; and the Unaligned and Reverse Format (URF), where the instance is represented as “$\mathtt{x_{0}\dots x_{n_{x}-1}+y_{0}\dots y_{n_{y}-1}=}$”, $n_{c}=\operatorname*{arg\,max}_{i}\{c_{i}\neq 0\},c\in\{x,y\}$. The formal scale in the language domain is the length of the string. However, the length of the input strings in neither format faithfully reflects the inherent problem scale: in ARF, the length of the input strings is always invariant; in URF, a shorter string may correspond to a larger scale than a longer string (e.g., $s_{1}=$$1+1234=$” is $4$-addition, $s_{2}=$$123+123=$” is $3$-addition; however, $|s_{1}|<|s_{2}|$). Furthermore, we find in experiments that Relative Position Embedding (RPE) improves the length generalization performance of Transformers for ARF, while it provides little benefit for the length generalization of Transformers for URF.
To develop robust and transferable methods for scalable models, it requires a formulation that captures the inherent scaling challenge of length generalization and is invariant to the nuisances such as data format. To do so, we first notice that in the length generalization of a wide range of problems, the instances can be seen as generated from some latent space whose dimension reflects the scale.
Example 2 (Addition).
Let $\Sigma=\{0,\dots,9\}\times\{0,\dots,9\}$. The instance of $n$-addition $x_{n-1}\dots x_{0}+y_{n-1}\dots y_{0}$ can be represented by the state vector $h_{n}=\left[\left(x_{0},y_{0}\right),\dots,\left(x_{n-1},y_{n-1}\right)\right]$. The $k$-th dimension of $h_{n}$ corresponds to the $(k-1)$-th digits of the two addenda (i.e., $x_{k-1}$ and $y_{k-1}$). The expansion of the state vector in the dimension corresponds to the increase in the addendum digits. Therefore, the length generalization from $N_{0}$-addition to $N$-addition can be seen as a generalization from the low-dimensional latent space $\Sigma^{N_{0}}$ to the high-dimensional latent space $\Sigma^{N}$.
To catch the above intuition, we make the following assumption about the underlying data generation.
Assumption 1 (Data Generation in Length Generalization).
The data generation process of an instance of scale $n$ with the concept $c$ is as follows:
  1. 1.
    A latent variable $h$ is sampled from a subspace $\Sigma^{n}$ of dimension $n$;
  2. 2.
    The label $y$ is determined by the concept $c$ and the latent variable $h$, i.e., $y=c(h)$;
  3. 3.
    The hidden variable $h$ is transformed to an input sequence by the data format mapping $\phi$.
According to Assumption 1, the inherent scaling challenge can be captured by Low-Dimension-to-High-Dimension (LDHD) generalization, a special scenario of OOD generalization with structural distributional shifts. In an LDHD generalization problem, a model is trained with samples from a low-dimensional subspace while evaluated with samples from a higher-dimensional superspace. To further illustrate, we can consider that the training samples are from the space $\Sigma^{N_{0}}$ and the testing samples are from $\Sigma^{N}$, where $\Sigma$ is a field and the numbers $N_{0},N$ satisfy $N_{0}<N$. With slight abuse of notation, we also use $\Sigma^{n}$ to denote the embedding space of the subspace $\Sigma$ into a space $\Sigma^{\bar{N}}$ for some sufficiently large $\bar{N}\geq N>N_{0}$ (we use $\bar{N}=N$ below if there is no special statement).
The main challenge of LDHD generalization is that the testing space has extra dimensions compared to the training space. As a result, the testing space contains instances with orthogonal components to the training space. The training samples can reveal no information about how these components contribute to the results. For instance, in Example 2, learning with $N_{0}$-addition solely does not tell how $(N_{0}+1)$-th digits to $N$-th digits contribute to the result unless we have the prior knowledge that the addition of each digit shares the same process. We formally this challenge as No-Free-Lunch Theorem of LDHD Generalization in Section 3.
The No-Free-Lunch Theorem [25, 24] of LDHD Generalization necessitates the use of prior knowledge in the learning process in order to achieve LDHD generalization. In practice, the prior knowledge is usually incorporated via the inductive bias of the learning algorithms and the models. Abbe et al. [1] claimed that models with (S)GD converge to min-degree interpolators. However, the minimal degree-profile bias neither holds uniformly for all model architectures nor is sufficient for achieving LDHD generalization. To develop LDHD generalizable models, we investigate different architectures whose inductive bias with (S)GD can potentially go beyond the min-degree interpolators. While random feature models are shown to be min-degree interpolators, we prove that random feature models with projections, where inputs are transformed by projections before fed to the random feature models, converge to min-degree interpolators under different independent sets. We further consider Position-Only Linear Attentions with Advice (PLAA), i.e., linear attentions whose attention scores are determined only by positions, with additional hints about the scales of the instances. This model can be seen as a simplified abstraction of decoder-only transformers with a special focus on positional relations that are considered crucial for length generalization. We demonstrate the failure of PLAA models equipped with Absolute Position Embeddings (APEs) in general cases and show the PLAA models with Relative Position Embeddings (RPEs) are min-degree interpolators w.r.t. some independent sets. These results show that for models biased towards min-degree interpolators w.r.t. some independent sets, LDHD generalization can be achieved if and only if the target mapping is the min-degree interpolator w.r.t. the same set.
The LDHD generalization perspective further provides insights on how Chain-of-Thought (CoT) [22] can help length generalization. We show that CoT can be seen as a change of the hidden space, where each dimension of the latent space is augmented with an additional middle state. This transformation could facilitate LDHD generalization in the latent space for some problems. Another implication is a principle of position embedding design for length generalization with Transformers: we need to handle both the inherent LDHD generalization and the nuisances such as the data format in the design of the position embeddings. Following this principle, we propose a novel position embedding named RPE-Square. The RPE-Square enhances the RPE with the ability to handle certain data format nuisances, evidently improving the length generalization of the URF addition over the Transformer with the RPE.
The main contributions of the paper are summarized as follows:
  • We propose to study LDHD generalization, which captures the inherent scaling challenge of a wide range of length generalization problems. LDHD generalization provides a formulation under which we can perform theoretical analysis for general length generalization problems.
  • We study the inductive bias of different model architectures under (S)GD, which can be exploited to incorporate prior knowledge to achieve LDHD generalization. In the context of Boolean functions, we show that different architectures under (S)GD can have inductive bias other than min-degree interpolators, which enables LDHD generalization for a wider class of problems.
Notations. We use $[n]$ to represent the set of numbers $\{1,\dots,N\}$. We denote the set of all functions from the set ${\mathcal{X}}$ to the set ${\mathcal{Y}}$ as ${\mathcal{F}}_{{\mathcal{X}},{\mathcal{Y}}}$. We define $\operatorname{Proj}(x,V)$ as the coordinate of the projection of $x$ onto the space spanned by $V$ with the basis $V=\left[v_{1},\dots,v_{r}\right]$, i.e., $\left[\operatorname{Proj}(x,V)\right]_{i}=\langle x,v_{i}\rangle$ for all $i=1,\dots,r$. We use $A^{*}$ to denote the Kleene closure of the set $A$, i.e., $A^{*}=\bigcup_{k=0}^{\infty}A^{k}$. We use $\deg(p)$ to denote the degree of the polynomial $p$.
2 Related Work
Length generalization in reasoning problems. Length generalization is a key challenge in learning to reason, typically interpreted as the ability to learn with small-scale instances of a task and generalize to unseen large instances of the same task [3, 31]. Various reasoning tasks are considered to investigate length generalization, including arithmetic [11, 7], Boolean logic [1, 6], symbolic reasoning [29], etc. Despite the rich literature on length generalization on specific reason tasks, few works have considered challenges and overconditions of length generalization for general problems. The existing works that analyze general length generalization mainly focus on the change of the input sequence length [26, 2], which does not precisely capture the difficulties of length generalization. Our work proposes to consider LDHD generalization, which characterizes the inherent challenges of length generalization in many tasks.
Inductive Bias of Model Architectures and Algorithms. In the situation of learning with overparatermization or incomplete information, proper inductive bias is essential to select the true model from the hypothesis [15, 5]. One of the most studied scenarios is the inductive bias of different model architectures under (S)GD. Previous research shows (S)GD combined with different model architectures lead to different effects of implicit regularization, such as different norms of the learnable parameters Gunasekar et al. [10], Bartlett et al. [5] and different complexity characterizations of the models [18, 19, 1]. We specially mention Abbe et al. [1], which proposes that models trained with (S)GD are biased towards min-degree profile interpolators in the context of Boolean functions, which do not achieve general LDHD generalization. Our results show that the min-degree profile bias does not hold for all models. We further show that models with different architecture can converge to min-degree profile interpolators under different independent sets when trained with (S)GD. This partially explains how model architectures can affect length generalization in reasoning problems.
3 Low-Dimension-to-High-Dimension Generalization
Definition 1 (Low-Dimension-to-High-Dimension Generalization).
Suppose that ${\mathcal{X}}$ is a sample space and ${\mathcal{X}}_{1},{\mathcal{X}}_{2}$ are two subspaces such that ${\mathcal{X}}_{1}\subset{\mathcal{X}}_{2}\subset{\mathcal{X}}$ and $\dim({\mathcal{X}}_{1})<\dim({\mathcal{X}}_{2})$. Consider a concept class ${\mathcal{C}}\subset{\mathcal{F}}_{{\mathcal{X}},{\mathcal{Y}}}$, two distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2}$ where $\operatorname{supp}({\mathcal{D}}_{1})={\mathcal{X}}_{1}$ and $\operatorname{supp}({\mathcal{D}}_{2})={\mathcal{X}}_{2}$, and a learning algorithm ${\mathcal{A}}:\left({\mathcal{X}}\times{\mathcal{Y}}\right)^{*}\mapsto{\mathcal{F}}_{{\mathcal{X}},{\mathcal{Y}}}$. We say low-dimension-to-high-dimension generalization of the concept class ${\mathcal{C}}$ from ${\mathcal{D}}_{1}$ to ${\mathcal{D}}_{2}$ is achieved by the algorithm ${\mathcal{A}}$ with $m$ samples and $\epsilon$ error if
$\mathbb{E}_{X^{m}\sim{\mathcal{D}}_{1}^{m},X_{m+1}\sim{\mathcal{D}}_{2}}\left[\ell\left(\hat{f}_{X^{m},c}(X_{m+1}),c(X_{m+1})\right)\right]\leq\epsilon,$ (1)
where $\hat{f}_{X^{m},c}$ is the function learned by the algorithm ${\mathcal{A}}$ from the training samples $X^{m}$ labeled by the concept $c$ and $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\mapsto{\mathbb{R}}$ is the loss function.
Definition 1 extends the Independent Identical Distribution (IID) assumption in PAC learning theory and considers a special shift between the training data and the testing data. This shift is particularly challenging because the testing space is of strictly higher dimension than the training space. Generally, it is impossible to fully capture the structure of the training space from the testing data as the training data reveal no information on how components in the orthogonal subspace contribute to the output. Therefore, there is no algorithm that can always guarantee to learn the concept from the training data. Theorem 1 formally states the nonexistence of universal algorithms for LDHD generalization.
Theorem 1 (No-Free-Lunch Theorem of LDHD Generalization).
Suppose that the two sets ${\mathcal{X}}$ and ${\mathcal{Y}}$ are finite. For some $N>N_{0}$, consider two subsets ${\mathcal{X}}_{N_{0}}$, ${\mathcal{X}}_{N}$ of ${\mathcal{X}}$ such that ${\mathcal{X}}_{N_{0}}\subsetneq{\mathcal{X}}_{N}\subseteq{\mathcal{X}}$ and $\dim({\mathcal{X}}_{N_{0}})=N_{0}<N=\dim({\mathcal{X}}_{N})$. Let $c_{1},c_{2}\in{\mathcal{F}}(:={\mathcal{F}}_{{\mathcal{X}}_{N},{\mathcal{Y}}})$ be two concepts such that $c_{1}(x)=c_{2}(x)$ for all $x\in{\mathcal{X}}_{N_{0}}$. For any $c\in{\mathcal{F}}$ and ${\mathcal{X}}^{\prime}\subseteq{\mathcal{X}}$, define ${\mathcal{F}}/\left(c\mid{\mathcal{X}}^{\prime}\right):=\left\{f\in{\mathcal{F}}\mid f(x)=c(x)\text{ for all }x\in{\mathcal{X}}^{\prime}\right\}$. Let $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\mapsto{\mathbb{R}}$ be the loss function. For any distribution ${\mathcal{D}}({\mathcal{X}}_{N})$ such that $\operatorname{supp}\left({\mathcal{D}}({\mathcal{X}}_{N})\right)={\mathcal{X}}_{N}$:
$\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{1}(x),f(x)\right)\right]=\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{2}(x),f(x)\right)\right].$
Theorem 1 necessitates the consideration of structural assumptions on the concept class such that a learning algorithm could identify the target concept from the hypothesis with the imperfect information provided by the low-dimensional training data. For example, the concept class of linear classifiers with fixed weight vector on ${\mathcal{X}}={\mathbb{R}}^{d}$, i.e., ${\mathcal{C}}=\{\operatorname{sgn}\left(w_{0}^{\intercal}x+b\right)\mid w_{0}\in{\mathbb{R}}^{d},b\in{\mathbb{R}}\}$ with the $d^{\prime}$-dimensional training sample space ${\mathcal{X}}_{1}={\mathbb{R}}^{d^{\prime}}\times\{0\}^{d-d^{\prime}}$ and the $d$-dimensional testing sample space where $d^{\prime}<d$. For any concept $c\in{\mathcal{C}}$, a learning algorithm could not identify the true concept $c$ solely from the training data from ${\mathcal{X}}_{1}$ and the hypothesis of all linear classifiers. However, if a linear algorithm exploits the structure of the concept class that the weight vector always equals $w_{0}$, it can compute the target bias from the training data in ${\mathcal{X}}_{1}$ and thus identify the target concept $c$.
We further investigate how model structures can affect LDHD generalization. We show in Section 4 that different models trained with (S)GD can be seen as min-degree interpolators under different functional bases in the context of Boolean functions, which is a joint inductive bias of the model structures and (S)GD. This suggests a principle for model design to achieve LDHD generalization: make the concept class “low-degree” under the independent set induced by the model structure.
Many generalization challenges in reasoning problems can be attributed to LDHD generalization inherently. Length generalization is representative of these challenges and of significance in reasoning problems. Intuitively, length generalization means that a model trained with small-scale instances of a reasoning problem can perform well on large-scale instances of the problem. To formalize the intuition, we consider a typical instance being generated from a hidden state variable $h$ that represents the core of the instance and is transformed to the model input by a data format mapping. The hidden state space is of the form $\Sigma^{*}$ for some field $\Sigma$. The dimension $n$ of the subspace from which the hidden state variable is sampled, $n=\operatorname*{arg\,max}_{k}\{h\in\Sigma^{k},h\not\in\Sigma^{k+1}\}$ , reflects the increase in the scale of the problem. Besides, we define the concept class on the hidden space, depicting that the change of the data format does not change the intrinsic logic of the concepts. Assumption 1 describes the pipeline of the data generation in length generalization.
4 Main Results
We show theoretically how different models succeed or fail to achieve LDHD generalization as the effect of the inductive bias of the architectures under (S)GD. We focus on Boolean functions. More specifically, in the context of Boolean functions, we have $\Sigma=\{\pm 1\}$, ${\mathcal{X}}=\Sigma^{N}$, ${\mathcal{X}}_{n}=\Sigma^{n}\times\{1\}^{N-n}$ for $n=1,\dots,N$. The set of all Boolean functions potentially considered is ${\mathcal{F}}={\mathcal{F}}_{{\mathcal{X}},{\mathbb{R}}}$. We consider LDHD generalization from $N_{0}$ to $N$ for some $N_{0}<N$. Define $I(f)$ as the minimal set $I$ of indices that the function $f$ can be represented as a function of $x_{I}$, i.e., $I(f):=\operatorname*{arg\,min}_{I\subset[N]}|I|$ such that $f(x)=\tilde{f}(x_{I})$ for some function $\tilde{f}$ and all $x\in{\mathcal{X}}$. We say a function $f$ is $k$-sparse if $|I(f)|\leq k$.
Before presenting the theoretical results, we introduce two concepts degree profile w.r.t. independent set and min-degree interpolator w.r.t. independent set, which extend the concept degree profile and the concept min-degree interpolator, respectively. We use the two concepts to characterize the inductive bias of different model architectures under (S)GD.
Definition 2 (Degree Profile w.r.t. Independent Set ${\mathcal{B}}$).
Suppose that =${\mathcal{B}}=\{b_{1},\dots,b_{R}\}$ is an independent set of functions in ${\mathcal{F}}$ and $D=\max_{b\in{\mathcal{B}}}\deg(b)$. Let $f\in{\mathcal{F}}$ be a function in the subspace spanned by ${\mathcal{B}}$, i.e., $f=\sum_{i=1}^{R}\hat{f}_{{\mathcal{B}}}(b_{i})b_{i}$ for some $\hat{f}_{{\mathcal{B}}}(b_{i})\in{\mathbb{R}},i=1,\dots,R$. The degree profile of the function w.r.t. ${\mathcal{B}}$, denoted by $\operatorname{DegP}_{{\mathcal{B}}}(f)$, is a $(D+1)$-tuple where $D_{i}=\sum_{b\in{\mathcal{B}},\deg(b)=D+1-i}\hat{f}_{{\mathcal{B}}}(b)^{2}$ for $i=1,\dots,D+1$. The order of degree profiles is identical to the lexicographic order of the corresponding $D$-tuples.
Definition 3 (Min-Degree Interpolator w.r.t. Independent Set ${\mathcal{B}}$).
Suppose that =${\mathcal{B}}=\{b_{1},\dots,b_{R}\}$ is an independent set of functions in ${\mathcal{F}}$. Let ${\mathcal{X}}^{\prime}$ be a subset of the sample space ${\mathcal{X}}=\{\pm 1\}^{N}$. Denote the set of all interpolators on ${\mathcal{X}}^{\prime}$ for the concept $c$ by ${\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$, i.e., ${\mathcal{G}}_{{\mathcal{X}}^{\prime},c}=\{g\in{\mathcal{F}}\mid g(x)=c(x)\text{ for all }x\in{\mathcal{X}}^{\prime}\}$. A function $g$ is called the min-degree interpolator w.r.t. ${\mathcal{B}}$ on ${\mathcal{X}}_{0}$ for the concept $c$ if $g\in{\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$ and $\operatorname{DegP}(g)\leq\operatorname{DegP}(g^{\prime})$ for all $g^{\prime}\in{\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$.
4.1 Random Feature Model with Projection
We first consider the random feature model (RFM) and a class of its variants, i.e., Random Feature Models with Projection (RFMP; see Definition 4). The original RFM can be seen as a special instance of RFMP with $V=I_{N}$.
Definition 4 (Random Feature Model with Projection).
Suppose that $V=[v_{1},\dots,v_{r}]\in{\mathbb{R}}^{N\times r}$ satisfies $V^{\intercal}V=I_{r}$. A random feature model with projection w.r.t. $V$ is
$f_{\text{RFMP}}^{V,K}(x;a)=\frac{1}{\sqrt{K}}\sum_{k=1}^{K}a_{k}\sigma\left(\left\langle w_{k},\operatorname{Proj}\left(x,V\right)\right\rangle+b_{k}\right),$
where $K$ is the number of random features, $a=\left[a_{1},\dots,a_{K}\right]^{\intercal}$ is the learnable parameter, $w_{k}\sim{\mathcal{N}}(0,I_{r}/r)$, $b_{k}\sim{\mathcal{N}}(0,1/r)$ for $k=1,\dots,K$, and $\sigma$ is the activation function.
Technically, we follow the strongly expressive condition [1] for the activation function $\sigma$. Abbe et al. [1] shows that the RFM converges to the min-degree interpolator when initialized at 00 and trained with GD. However, this is not the case for all RFMP models. We show in Theorem 2 that an RFMP model converges to a min-degree interpolator w.r.t. an independent set determined by $V$.
Theorem 2.
Suppose that $V=[v_{1},\dots,v_{r}]\in{\mathbb{R}}^{N\times r}$ satisfies $V^{\intercal}V=I_{r}$. Define the set ${\mathcal{B}}(V)$ of independent functions as
${\mathcal{B}}(V)=\left\{\chi^{V}_{T}(x)\right\}_{T\subseteq[r]},\quad\text{where }\chi^{V}_{T}(x)=\prod_{t\in T}\sum_{n=1}^{N}(v_{t})_{n}x_{n}$
Let $a_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $a$ is initialized at $a_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},c,V}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c^{*}(x)=c\left(\operatorname{Proj}(x,V)\right)$ that is $O_{N}(1)$-sparse. Then we have
$f_{\text{RFMP}}^{V,K}(x;a_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},c,V}}\operatorname{DegP}_{{\mathcal{B}}(V)}(g),\quad\text{as }K\to\infty,t\to\infty.$
When $V=I_{N}$, the independent set ${\mathcal{B}}(V)$ is the Fourier basis of the Boolean functions, and Theorem 2 implies that the RFM converges to the min-degree interpolator. From Theorem 2, we see that an RFMP model with the projection matrix $V$ can achieve LDHD generalization only if the target concept coincides with the min-degree interpolator w.r.t. ${\mathcal{B}}(V)$. Specially, for the RFM, we have:
Corollary 1.
For any $f\in{\mathcal{F}}$ such that $I(f)\not\subset\left[N_{0}\right]$, the min-degree interpolator does not achieve LDHD generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ and thus the RFM initialized at 0 and trained with GD does not achieve LDHD generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$.
Corollary 1 shows that the min-degree interpolator and thus the RFM model can only achieve LDHD generalization for a very restricted set of functions that are only dependent on $x_{[N_{0}]}$. To achieve LDHD generalization with RFMP models requires prior knowledge of the concept class to design a projection to align the concepts with the min-degree interpolator w.r.t. the projected Fourier basis.
4.2 Position-Only Linear Attention with Advice
In this subsection, we investigate Position-Only Linear Attention with Advice (PLAA), which can be seen as a simplification of decoder-only Transformers (Definition 5), with a special focus position embeddings that are considered pivot to the length generalization of the Transformers [20, 11].
Definition 5 (PLAA).
Define the advice function $n:{\mathcal{X}}\mapsto\{0,\dots,N\}$ such that $n(x)=\operatorname*{arg\,max}_{n}\{x_{n}=-1\}$ if there exists $k\in[N]$ such that $x_{k}=-1$ and $n(x)=0$ otherwise. We additionally define $e_{0}:=0$. A PLAA model is
$f_{\text{PLAA}}(x;A)=x^{\intercal}Ae_{n(x)},$
where $A\in{\mathcal{U}}_{N}$ is the learnable parameter and $e_{n}$ denotes the vector with a $1$ in the $n$-th coordinate and 00’s elsewhere.
We further elaborate on the intuition behind the PLAA models. In the generation process of a decoder-only (linear) Transformer with position embeddings given input $\mathtt{s=s_{1}\dots s_{n}}$, the attention is computed by the query at the position $n$ and the keys at the positions $i\leq n$. The position of the query is special, advising the length of the input and reflecting the scale of the instance ideally. The PLAA model captures this feature and introduces the notation $n(x)$ to reflect the dimension of the subspace that $x$ belongs to. To further simplify and focus on the position embeddings, we assume that the value of each $x_{i}$ is identical to itself and the attention is only related to positions. The contribution of the interaction between the position embeddings is $\left[e_{1},\dots,e_{n}\right]^{\intercal}A_{[n],[n]}e_{n}=A_{[n],n}$ for some upper triangle matrix $A$. When the input is embedded to length $N$ but the query is still made at the position $n$, the output of the model is $x^{\intercal}Ae_{n(x)}$, i.e., the output of the PLAA model. Therefore, the PLAA model is a simplification of decoder-only Transformers that enables theoretical analysis of the impact of the position embeddings on length generalization.
In Definition 5, we directly parameterize the PLAA model directly with the attention matrix. In practice, however, the attention matrix is typically computed by the interaction between the position embeddings. Therefore, we consider the PLAA models with the Absolute Position Embedding (APE) and the Relative Position Embedding (RPE), respectively. See Definitions 6 and 7. Note that we consider Generalized RPE (GPRE) in Definition 7 because the RPE can be seen as a special instance of the GRPE with ${\mathcal{U}}={\mathcal{U}}_{\text{RPE}}=\{D_{1},\dots,D_{N}\}$, where $D_{k}$ is a $k$-th upper diagonal matrix such that $(D_{k})_{ij}$ is $1$ if $j=i+k-1$ and 00 otherwise. We seek a more general result applicable to all similar parameterization methods to the RPE.
Definition 6 (PLAA with APE).
A PLAA model with APE is
$f_{\text{PLAA}}^{\text{APE}}(x;P)=x^{\intercal}\left(M_{N}^{u}\circ P^{\intercal}P\right)e_{n(x)},$
where $M_{N}^{u}{\mathbb{R}}^{N\times N}$ is the upper triangle mask, i.e., $(M_{N}^{u})_{ij}$ is $1$ if $i\leq j$ and 00 otherwise, and $P\in{\mathbb{R}}^{d_{P}\times N}$ is the learnable parameter of the model.
Definition 7 (PLAA with GRPE).
For ${\mathcal{U}}=\{U_{1},\dots,U_{r}\}$, a PLAA moodel with GRPE is
$f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p)=x^{\intercal}\left(\sum_{i=1}^{r}U_{i}p_{i}\right)e_{n(x)},$
where $U_{i}\in{\mathcal{U}}_{N},i=1,\dots,r$ are upper triangle matrices that satisfy $\left\langle U_{i},U_{i}\right\rangle=1$ for all $i=1,\dots,r$ and $\left(U_{i}\right)_{kl}\left(U_{j}\right)_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$, and $p=\left[p_{1},\dots,p_{r}\right]^{\intercal}\in{\mathbb{R}}^{r}$ is the learnable parameter of the model.
Remark 1.
The condition $\left(U_{i}\right)_{kl}\left(U_{j}\right)_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$ in Definition 7 means each element in the position-only attention is characterized by at most one parameter. For example, in RPE, $A_{i,j}(i\leq j)$ is only parameterized by $p_{j-i}$. This condition naturally generalizes the intuition of RPE that characterizes some “relative” relations between two positions.
We first consider the vanilla PLAA model, the direct parameterization with the attention matrix. Theorem 3 shows that the PLAA model converges to the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{N}^{\text{PLAA}}$ that is defined in the theorem.
Theorem 3.
Define the set ${\mathcal{B}}_{N}^{\text{PLAA}}$ as
${\mathcal{B}}_{N}^{\text{PLAA}}:=\{b_{ij}^{\text{PLAA}}(x)\}_{1\leq i\leq j\leq N},$
where
$b_{ij}^{\text{PLAA}}(x)=\begin{cases}-\frac{1-x_{j}}{2}\prod_{k=j+1}^{N}\frac{1+x_{k}}{2},&i=j,\\ x_{i}\frac{1-x_{j}}{2}\prod_{k=j+1}^{N}\frac{1+x_{k}}{2},&i<j.\end{cases}$
Let $A_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $A$ is initialized at $A_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}(x;A^{*})$. Then we have
$f_{\text{PLAA}}(x;A_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}}\operatorname{DegP}_{{\mathcal{B}}_{N}^{\text{PLAA}}}(g),\quad\text{as }t\to\infty.$
For the PLAA model with APE, Theorem 4 shows that the model converges to the interpolator whose attention matrix is of the minimal nuclear norm if the minimal-nuclear-norm-solution conjecture (Assumption 2 about the implicit regularization in matrix factorization holds. While the conjecture has not been theoretically justified in general cases, it is proved correct under some specific conditions and supported by empirical evidence [10, 4].
Assumption 2 (Minimal-Nuclear-Norm-Solution Conjecture [10]).
Consider the optimization
$\min_{U\in{\mathbb{R}}^{n\times n}}f(U)=\left\|{\mathcal{A}}\left(UU^{\intercal}\right)-y\right\|_{2}^{2},$
where ${\mathcal{A}}:{\mathbb{R}}^{n\times n}\mapsto{\mathbb{R}}^{m}$ is a linear operator specified by ${\mathcal{A}}(\cdot)_{i}=\left\langle A_{i},\cdot\right\rangle$ for symmetric linearly independent $A_{i}\in{\mathbb{R}}^{n\times n},1\leq i\leq m$, and $y\in{\mathbb{R}}^{m}$. Define $X=X(U):=UU^{\intercal}$. Let $U_{t}$ be the parameter reached by gradient flow in the timestep $t$ and $X_{t}$ be $U_{t}U_{t}^{\intercal}$. Define $X_{\infty}(X_{\text{init}}):=\lim_{t\to\infty}X_{t}$ for $U_{0}U_{0}^{\intercal}=X_{\text{init}}$. For any full rank $X_{0}$, if $\hat{X}=\lim_{\alpha\to 0}X_{\infty}(\alpha X_{0})$ exists and ${\mathcal{A}}\left(\hat{X}\right)=y$, then $\hat{X}\in\operatorname*{arg\,min}_{{\mathcal{X}}\succ\|X\|_{*}}$ s.t. ${\mathcal{A}}(X)=y$.
Theorem 4.
Suppose that Assumption 2 holds and $d_{P}=N$. Let $P_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $P$ is initialized at $P_{0}$ such that $(P_{0})_{ij}\sim{\mathcal{N}}(0,\alpha)$ for $(i,j)\in[d_{P}]\times[N]$ and $\alpha$ sufficiently small, and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},P^{*}}^{\text{APE}}$ be the set of all PLAA interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}^{\text{APE}}(x;P^{*})$. Then we have
$f_{\text{PLAA}}^{\text{APE}}(x;P_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},P^{*}}^{\text{APE}}}\|A_{g}\|_{*},\quad\text{as }t\to\infty,$
where $A_{g}$ is the attention matrix in the PLAA model $g$.
Theorem 4 shows that APE leads to a low-rank attention matrix in the PLAA model. Therefore, when the training data is restricted to ${\mathcal{X}}_{N_{0}}$, the rank of the attention matrix learned by the model with APE is no higher than $N_{0}$. This leads to length generalization failure when the target concept requires higher-rank attention to be represented. We have the following direct corollary of Theorem 4 that partially explains the limitation of APE for length generalization.
Corollary 2.
The PLAA with APE fails to achieve LDHD generalization for the concept $c(x)=f_{\text{PLAA}}(x;A^{*})$ if $\operatorname{rank}(A^{*})>N_{0}$.
Theorem 5 characterizes that the PLAA converges to the interpolator that minimizes the degree-profile w.r.t. to the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$, which can be seen as the projections of ${\mathcal{U}}$ onto ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$. Theorem 3 is a special case of Theorem 5 for ${\mathcal{U}}=\left\{E_{ij}\right\}_{1\leq i\leq j\leq N}$, where $E_{ij}\in{\mathbb{R}}^{N\times N}$ is the elementary matrix that has a 1 at the $(i,j)$-th position and 0 everywhere else.
Theorem 5.
For the ${\mathcal{U}}=\{U_{1},\dots,U_{r}\}$, define
${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}:=\left\{\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)\right\}_{1\leq k\leq r}.$
Let $p_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $p$ is initialized at $p_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},p^{*}}^{\text{GRPE},{\mathcal{U}}}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p^{*})$. Then we have
$f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},p^{*}}^{\text{GRPE},{\mathcal{U}}}}\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(g),\quad\text{as }t\to\infty.$
With the inductive bias of the PLAA with GRPE, Corollary 3 states that LDHD generalization can be achieved if and only if the target concept can be represented by the elements in ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ that have dependence on ${\mathcal{X}}_{N_{0}}$.
Corollary 3.
Consider the concept $c(x)=f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p^{*})=\sum_{k=1}^{r}c_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$. Under the conditions of Theorem 5, the PLAA with GRPE achieves LDHD generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ if and only if
$\left\{k\mid(U_{k})_{[N_{0}],[N_{0}]}=0\right\}\subseteq\left\{k\mid c_{k}=0\right\}.$
Remark 2.
While the projection in the RFM and the position embeddings in the PLAA may introduce a stronger inductive bias that benefits LDHD generalization, they can also reduce the expressiveness of the models. The critical point is to align the models with the concept class in such a way that we can incorporate a strong inductive bias for LDHD generalization while maintaining sufficient expressiveness for the concept class.
5 Further Implications for Length Generalization
We discuss further implications of the LDHD generalization perspective for length generalization.
5.1 Chain-of-Thought for Length Generalization
While the Chain-of-Thought (CoT) can lead to more variety in the length of testing samples, it is widely used as an effective technique to improve the length generalization in various reasoning tasks. This seems contradictory if considered in the original space of the input sequence. However, when considering the abstraction of the data generating process from the hidden space, we can see that CoT intrinsically changes the underlying hidden space by extending each dimension with a “middle” variable and does not lead to the dimensional increase in the hidden space. For example, consider the $n$-addition without CoT and the $n$-addition with CoT. In the case without CoT, the instance $x_{n-1}\dots x_{0}+y_{n-1}\dots y_{0}=z_{n}\dots z_{0}$ corresponds to the latent state $h_{n}=\left[\left(x_{0},y_{0}\right),\dots,(x_{n-1},y_{n-1})\right]\in\Sigma^{n}$ for $\Sigma=\{0,\dots,9\}\times\{0,\dots,9\}$. In the case with CoT, one step of predicting $z_{t}$ in solving the instance corresponds to the latent state $h_{n}=\left[\left(x_{0},y_{0},z_{0}\right),\dots,\left(x_{t-1},y_{t-1},z_{t-1}\right),\left(x_{t},y_{t},*\right)\dots,(x_{n-1},y_{n-1},*)\right]\in\bar{\Sigma}^{n}$ for $\bar{\Sigma}=\Sigma\times\{*,0,\dots,9\}$, where * is a special element indicating undetermined values. CoT does not cause the LDHD generalization challenge but extends the field $\Sigma$, which could potentially lead to a more easily learnable target concept under the model architecture.
5.2 Position Embeddings for Length Generalization
Position embeddings are considered closely related to the length generalization in Transformers Our analysis suggests a principle to consider both the inherent LDHD generalization and the nuisances such as the data format in the design of the position embeddings. To further elaborate, consider the length generalization of the URF addition with CoT. While RPE can capture the recursive structure of the addition problem and could lead to LDHD generalization in the latent space, it fails to work for the URF addition with CoT. The failure of RPE is attributed to the nuisance of the URF.
[…]
$\sum_{1\leq k\leq j,1\leq l\leq i}\frac{\exp\left((W_{Q}x_{j})^{\intercal}(W_{K}x_{l})\right)}{\sum_{1\leq l^{\prime}\leq j}\exp\left((W_{Q}x_{j})^{\intercal}(W_{K}x_{l^{\prime}})\right)}\frac{\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k})\right)}{\sum_{1\leq k^{\prime}\leq i}\exp\left((W_{Q}x_{i})^{\intercal}(W_{K}x_{k^{\prime}})\right)}R_{(j-l)-(i-k)},$
[…]
Figure 1: Length generalization of Transformer with RPE and RPE-Square in the URF addition. The training data are from the URF $4$-addition and the testing data are from the URF $5$-addition. (a) Both the models are trained for $20000$ steps. The comparison result shows that the RPE fails while the RPE-Square succeeds in achieving length generalization in the URF addition, (b) The RPE-Square model is trained for $200000$ steps, which shows the RPE-Square model converges to a solution achieving nearly perfect accuracy. The “$k$-th” line represents the digitwise accuracy to predict the $k$-th digit, i.e. $z_{k}$.
6 Conclusion and Discussion
We propose considering LDHD generalization, which captures the inherent scaling challenge of length generalization in various reasoning tasks. We introduce the No-Free-Lunch Theorem for LDHD generalization, demonstrating the necessity of inductive bias for achieving length generalization. We further investigate the inductive biases of different model architectures trained with (S)GD in the context of Boolean functions. LDHD generalization can only be achieved if the inductive bias aligns with the structure of the target concept. As implications for length generalization in practical problems, our perspective on LDHD generalization elucidates the role of CoT in extending the latent space and leads to the principle that both inherent LDHD generalization and nuisances must be addressed to achieve length generalization.
For future work, while our theory is established with simplified models, it is crucial to further investigate inductive bias in more practical and complex models. Additionally, developing a paradigm for position embedding design based on the proposed principles is a valuable avenue for exploration.
Appendix A Proofs
A.1 Proof for Theorem 1
Let ${\mathcal{D}}_{1}:={\mathcal{D}}\left({\mathcal{X}}_{N}\mid x\in{\mathcal{X}}_{N_{0}}\right)$ be the conditional distribution given that $x\in{\mathcal{X}}_{N_{0}}$ and ${\mathcal{D}}_{2}:={\mathcal{D}}\left({\mathcal{X}}_{N}\mid x\not\in{\mathcal{X}}_{N_{0}}\right)$ be the conditional distribution given that $x\not\in{\mathcal{X}}_{N_{0}}$
$\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{1}(x),f(x)\right)\right]$
$=$ $P_{{\mathcal{D}}({\mathcal{X}}_{N})}(x\in{\mathcal{X}}_{N_{0}})\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{1}(x),f(x)\right)\right]$
$+P_{{\mathcal{D}}({\mathcal{X}}_{N})}(x\not\in{\mathcal{X}}_{N_{0}})\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{1}(x),f(x)\right)\right]$
$=$ $P_{{\mathcal{D}}({\mathcal{X}}_{N})}(x\in{\mathcal{X}}_{N_{0}})V_{1}(c_{1})+P_{{\mathcal{D}}({\mathcal{X}}_{N})}(x\not\in{\mathcal{X}}_{N_{0}})V_{2}(c_{1}),$
where
$V_{1}(c):=\sum_{f\in{\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c(x),f(x)\right)\right],$
$V_{2}(c):=\sum_{f\in{\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c(x),f(x)\right)\right].$
Similarly, we have
$\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{2}(x),f(x)\right)\right]=P_{{\mathcal{D}}({\mathcal{X}}_{N})}(x\in{\mathcal{X}}_{N_{0}})V_{1}(c_{2})+P_{{\mathcal{D}}({\mathcal{X}}_{N})}(x\not\in{\mathcal{X}}_{N_{0}})V_{2}(c_{2}).$
To prove the theorem, it remains to show that $V_{1}(c_{1})=V_{1}(c_{2})$ and $V_{2}(c_{1})=V_{2}(c_{2})$.
By the definition of ${\mathcal{F}}/\left(c|_{{\mathcal{X}}_{N_{0}}}\right)$, we have
$V_{1}(c_{1})$ $=\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{1}(x),f(x)\right)\right]$
$=\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{1}(x),c_{1}(x)\right)\right]$
=$\overset{(a)}{=}\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{2}(x),c_{2}(x)\right)\right]$
$=\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{1}}\left[\ell\left(c_{2}(x),f(x)\right)\right]=V_{1}(c_{2}),$
where the equality (a) is due to that $c_{1}(x)=c_{2}(x)$ for all $x\in{\mathcal{X}}_{N_{0}}$.
By the no-free-lunch theorem [25, 24] for ${\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}}$, ${\mathcal{Y}}$, and ${\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}$, we have
$\sum_{f\in{\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{1}(x),f(x)\right)\right]=\sum_{f\in{\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{2}(x),f(x)\right)\right].$
Then we have
$V_{2}(c_{1})$ $=\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{1}(x),f(x)\right)\right]$
$=\sum_{f\in{\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{1}(x),f(x)\right)\right]$
$=\sum_{f\in{\mathcal{F}}_{{\mathcal{X}}_{N}\setminus{\mathcal{X}}_{N_{0}},{\mathcal{Y}}}}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{2}(x),f(x)\right)\right]$
$=\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}_{2}}\left[\ell\left(c_{2}(x),f(x)\right)\right]=V_{2}(c_{2}).$
A.2 Proof for Theorem 2
Let $z=z(x)=\operatorname{Proj}(x,V)$ and ${\mathcal{Z}}_{N_{0}}=\left\{\operatorname{Proj}(x,V)\mid x\in{\mathcal{X}}_{N_{0}}\right\}$. Let ${\mathcal{G}}^{\prime}_{N_{0},c}$ be the set of all RFM interpolators on ${\mathcal{Z}}_{N_{0}}$ for the concept $c(z)$, i.e.,
${\mathcal{G}}^{\prime}_{N_{0},c}=\{f_{\text{RFM}}(z;a)\mid f_{\text{RFM}}(z;a)=c(z)\text{ for all }z\in{\mathcal{Z}}_{N_{0}}\},$
where
$f_{\text{RFM}}(z;a)=\frac{1}{\sqrt{K}}\sum_{k=1}^{K}a_{k}\sigma(\langle w_{k},z\rangle+b_{k}).$
According to the proof of Theorem 3.8 in Abbe et al. [1], the RFM model $f_{\text{RFM}}(z;a)$ converges to the min-degree interpolator (w.r.t. the variable $z$) in ${\mathcal{G}}^{\prime}_{N_{0},c}$, i.e.,
$f_{\text{RFM}}(z;a_{t})\to\operatorname*{arg\,min}_{g^{\prime}\in{\mathcal{G}}^{\prime}_{N_{0},c}}\operatorname{DegP}_{B^{\prime}}(g^{\prime}),\quad\text{as}K\to\infty,t\to\infty,$
where $B^{\prime}=\{\chi_{T}(z)\}_{T\subseteq[r]}$.
By the definition of $z=z(x)=\operatorname{Proj}(x,V)$, we have
$f_{\text{RFMP}}^{V,K}(x;a_{t})=f_{\text{RFM}}(z;a_{t})\to\operatorname*{arg\,min}_{g^{\prime}\in{\mathcal{G}}^{\prime}_{N_{0},c}}\operatorname{DegP}_{B^{\prime}}(g^{\prime})=\arg\min_{g\in{\mathcal{G}}_{N_{0},c,V}}\operatorname{DegP}_{{\mathcal{B}}(V)}(g),$
as $K\to\infty,t\to\infty$.
A.3 Proof for Theorem 5
Since Theorem 3 is a special case of Theorem 5 for ${\mathcal{U}}=\{E_{ij}\}_{1\leq i\leq j\leq N}$, we only need to prove Theorem 5.
Note that
$f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p)=x^{\intercal}\left(\sum_{k=1}^{r}p_{k}U_{k}\right)e_{n(x)}=\left\langle xe_{n(x)}^{\intercal},\sum_{k=1}^{r}p_{k}U_{k}\right\rangle.$
Let $D_{N_{0}}$ be the size of the training set in ${\mathcal{X}}_{N_{0}}$. Then the loss can be represented as
$L(p)=\frac{1}{2D_{N_{0}}}\sum_{x\in{\mathcal{X}}_{N_{0}}}\left(\left\langle xe_{n(x)}^{\intercal},\sum_{k=1}^{r}p_{k}U_{k}\right\rangle-c^{*}(x)\right)^{2}.$
Without loss of generality, we use the notation of gradient flow in this proof. Then we have
$\dot{A}=-\frac{1}{D_{N_{0}}}\sum_{x\in{\mathcal{X}}_{N}}\left(\left\langle xe_{n(x)}^{\intercal},A\right\rangle-c^{*}(x)\right)\sum_{k=1}^{r}\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle U_{k}.$
Since $p(0)=0$, we have $A(0)=0$ and thus $A(t)\in\operatorname{span}\left\{\sum_{k=1}^{r}\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle U_{k}\right\}_{x\in{\mathcal{X}}_{N_{0}}}$. Then the convergence point $\hat{A}$ can be represented as $\hat{A}=\sum_{x\in{\mathcal{X}}_{N_{0}}}\hat{a}(x)\sum_{k=1}^{r}\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle U_{k}$. The convergence function $\hat{f}$ is
$\hat{f}(x)$ $=\sum_{x^{\prime}\in{\mathcal{X}}_{N_{0}}}\hat{a}(x^{\prime})\sum_{k=1}^{r}\left\langle x^{\prime}e_{n(x^{\prime})}^{\intercal},U_{k}\right\rangle\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle$
$=\sum_{k=1}^{r}\sum_{x^{\prime}\in{\mathcal{X}}_{N_{0}}}\hat{a}(x^{\prime})\left\langle x^{\prime}e_{n(x^{\prime})}^{\intercal},U_{k}\right\rangle\left\langle xe_{n(x)}^{\intercal},U_{k}\right\rangle$
Lemma 1.
For any upper triangle matrix $U\in{\mathbb{R}}^{N\times N}$, we have
$\langle xe_{n(x)}^{\intercal},U\rangle=\sum_{1\leq i\leq j\leq N}U_{ij}b_{ij}^{\text{PLAA}}(x).$
Proof for Lemma 1.
We first prove that $\langle xe_{n(x)}^{\intercal},E_{mn}\rangle=b_{ij}^{\text{PLAA}}(x)$ for any $1\leq i\leq j\leq N$. Notice that
$\langle xe_{n(x)}^{\intercal},E_{ij}\rangle=\begin{cases}x_{i}&x_{j}=-1\land x_{j+1}=1\land\dots\land x_{N}=1,\\ 0&\text{otherwise}.\end{cases}$
Therefore, we have
$\langle xe_{n(x)}^{\intercal},E_{ij}\rangle$ $=x_{i}\cdot\frac{1-x_{j}}{2}\cdot\frac{1+x_{j+1}}{2}\cdot\dots\cdot\frac{1+x_{N}}{2}$
=$=\begin{cases}-\frac{1-x_{j}}{2}\prod_{k=j+1}^{N}\frac{1+x_{k}}{2},&i=j,\\ x_{i}\frac{1-x_{j}}{2}\prod_{k=j+1}^{N}\frac{1+x_{k}}{2},&i<j.\end{cases}$
$=b_{ij}^{\text{PLAA}}(x).$
For any upper triangle matrix $U=\sum_{1\leq i\leq j\leq N}U_{ij}E_{ij}$, we have
$\langle xe_{n(x)}^{\intercal},U\rangle=\langle xe_{n(x)}^{\intercal},\sum_{1\leq i\leq j\leq N}U_{ij}E_{ij}\rangle=\sum_{1\leq i\leq j\leq N}U_{ij}\langle xe_{n(x)}^{\intercal},E_{ij}\rangle=\sum_{1\leq i\leq j\leq N}U_{ij}b_{ij}^{\text{PLAA}}(x).$
Define $b_{k}(x):=\langle xe_{n(x)}^{\intercal},U_{k}\rangle$. By Lemma 1, we have $b_{k}(x)=\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$. Then the convergence function $\hat{f}$ can be represented as
$\hat{f}(x)=\sum_{k=1}^{r}\sum_{x^{\prime}\in{\mathcal{X}}_{N_{0}}}\hat{a}(x^{\prime})b_{k}(x^{\prime})b_{k}(x)=\sum_{k=1}^{r}\hat{p}_{k}b_{k}(x),$
where $\hat{p}_{k}=\sum_{x^{\prime}\in{\mathcal{X}}_{N_{0}}}\hat{a}(x^{\prime})b_{k}(x^{\prime})$.
For any interpolator $g(x)=\sum_{k=1}^{r}p_{k}b_{k}(x)$ on ${\mathcal{X}}_{N_{0}}$:
  • If $b_{k}(x)=0$ for all $x\in{\mathcal{X}}_{0}$, we have $\hat{p}_{k}=0$ and $p_{k}^{2}\geq 0=\hat{p}_{k}^{2}$;
  • If $b_{k}(x)\neq 0$ for some $x\in{\mathcal{X}}_{0}$, we have $\hat{p}_{k}=p_{k}$ and thus $\hat{p}_{k}^{2}=p_{k}^{2}$. To show this, without loss of generality, we suppose that for $1\leq k\leq m$, there exists some $x\in{\mathcal{X}}_{N_{0}}$ such that $b_{k}(x)\neq 0$, and for $m+1\leq k\leq r$, $b_{k}(x)=0$ for all $x\in{\mathcal{X}}_{N_{0}}$. It suffices to show that the following equation has a unique solution:
    $\begin{bmatrix}b_{1}(x_{1})&\dots&b_{m}(x_{1})\\ \vdots&\vdots&\vdots\\ b_{1}(x_{M})&\dots&b_{m}(x_{M})\end{bmatrix}\begin{bmatrix}p_{1}\\ \vdots\\ p_{m}\end{bmatrix}=\begin{bmatrix}c^{*}(x_{1})\\ \vdots\\ c^{*}(x_{M})\end{bmatrix},$
    where $M=|{\mathcal{X}}_{N_{0}}|$ and ${\mathcal{X}}_{N_{0}}=\{x_{1},\dots,x_{M}\}$. Since $p_{1}=\hat{p}_{1},\dots,p_{m}=\hat{p}_{m}$ is a solution, it remains to prove the uniqueness. Let
    $B_{m}=\begin{bmatrix}b_{1}(x_{1})&\dots&b_{m}(x_{1})\\ \vdots&\vdots&\vdots\\ b_{1}(x_{M})&\dots&b_{m}(x_{M})\end{bmatrix}.$
    It suffices to show $\operatorname{rank}(B_{m})=m$. Note that $\left(B_{m}^{\intercal}B_{m}\right)_{ii}=\sum_{x\in{\mathcal{X}}_{N_{0}}}b_{i}(x)^{2}>0$ and $\left(B_{m}^{\intercal}B_{m}\right)_{ij}=\sum_{x\in{\mathcal{X}}_{N_{0}}}b_{i}(x)b_{j}(x)=0$. We have $\operatorname{rank}(B_{m}^{\intercal}B_{m})=m$ and thus $\operatorname{rank}(B_{m})=m$.
Hence, for any interpolator $g$ on ${\mathcal{X}}_{N_{0}}$, we have
$\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(\hat{f})\leq\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(g),$
or equivalently,
$f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},p^{*}}^{\text{GRPE},{\mathcal{U}}}}\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(g),\quad\text{as }t\to\infty.$
A.4 Proof for Corollary 3
By Theorem 5, the PLAA with GRPE achieves LHDH generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ if and only if the target concept $c(x)$ is the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$. Therefore, it is equivalent to prove the target concept $c(x)$ is the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ if and only if
$\left\{k\mid(U_{k})_{[N_{0}],[N_{0}]}=0\right\}\subseteq\left\{k\mid c_{k}=0\right\}.$
We denote $\left\{k\mid(U_{k})_{[N_{0}],[N_{0}]}=0\right\}$ by ${\mathcal{K}}_{1}$ and $\left\{k\mid c_{k}=0\right\}$ by ${\mathcal{K}}_{2}$ in this proof.
We first show the sufficiency. Assume that $c(x)$ is not the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ on ${\mathcal{X}}_{N_{0}}$. In other words, there exists an interpolator $\tilde{c}(x)=\sum_{k=1}^{r}\tilde{c}_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$ on ${\mathcal{X}}_{N_{0}}$ such that
$\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(\tilde{c})<\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(c).$
Since both $c(x)$ and $\tilde{c}(x)$ are interpolators on ${\mathcal{X}}_{N_{0}}$, we have
$\sum_{k=1}^{r}c_{k}(U_{k})_{[N_{0}],[N_{0}]}=\sum_{k=1}^{r}\tilde{c}_{k}(U_{k})_{[N_{0}],[N_{0}]},$
or equivalently,
$\sum_{k\not\in{\mathcal{K}}_{1}}c_{k}(U_{k})_{[N_{0}],[N_{0}]}=\sum_{k\not\in{\mathcal{K}}_{1}}\tilde{c}_{k}(U_{k})_{[N_{0}],[N_{0}]}.$
Since $(U_{i})_{kl}(U_{j})_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$, we have $\left\{(U_{k})_{[N_{0}],[N_{0}]}\right\}_{k\not\in{\mathcal{K}}}$ are linearly independent. This implies $c_{k}=\tilde{c}_{k}$ for all $k\not\in{\mathcal{K}}_{1}$.
Since ${\mathcal{K}}_{1}\subseteq{\mathcal{K}}_{2}$, we have $c_{k}=0$ and thus $c_{k}^{2}\leq\tilde{c}_{k}^{2}$ for all $k\in{\mathcal{K}}_{1}$. Hence, we have
$\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(c)\leq\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(\tilde{c}),$
which contradicts the assumption.
We then prove the necessity. Assume that ${\mathcal{K}}_{1}\not\subseteq{\mathcal{K}}_{2}$. Then there exists some $k_{0}\in{\mathcal{K}}_{1}$ but $k_{0}\not\in{\mathcal{K}}_{2}$. Define $\tilde{\tilde{c}}(x)=\sum_{k=1}^{r}\tilde{\tilde{c}}_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$ where
$\tilde{\tilde{c}}_{k}=\begin{cases}0,&k=k_{0},\\ c_{k},&k\neq k_{0}.\\ \end{cases}$
By the definition of ${\mathcal{K}}_{1}$, $\tilde{\tilde{c}}(x)=c(x)$ for all $x\in{\mathcal{X}}_{N_{0}}$ and thus $\tilde{\tilde{c}}(x)$ is also an interpolator on ${\mathcal{X}}_{0}$. Note that $c_{k_{0}}\neq 0$. By the definition of $\tilde{\tilde{c}}(x)$, we have
$\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(\tilde{\tilde{c}})\leq\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(c),$
which contradicts that the target concept $c(x)$ is the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ on ${\mathcal{X}}_{N_{0}}$.
Appendix B Experiments
B.1 Position Embeddings for Length Generalization
For the URF $n$-addition training data, we first sample the lengths of two addends uniformly from $\{1,\dots,n\}\times\{1,\dots,n\}$. For two addends of lengths $(n_{1},n_{2})$, we then samples from $[10^{n_{1}-1},10^{n_{1}}-1]\times[10^{n_{2}-1},10^{n_{2}}-1]$ (If $n_{i}=0$, then the corresponding sample interval is $[0,9]$). This is to guarantee the addends are length-uniform. For the addition $x_{n_{1}-1}\dots x_{0}+y_{n_{2}-1}\dots y_{0}=z_{n_{3}}z_{n_{3}-1}\dots z_{0}$, the training instance is
$\mathtt{b\,x_{0}\,\dots\,x_{n_{1}-1}\,+\,y_{0}\,\dots\,y_{n_{2}-1}\,=\,z_{0}\,\dots\,z_{n_{3}-1}\,z_{n_{3}}\,e}.$
Here, we add spaces between the characters to ensure each is tokenized separately. We use “$\mathtt{b}$” and “$\mathtt{e}$” instead of “$\mathtt{[BOS]}$” and “$\mathtt{[EOS]}$” for a simpler implementation with the GPT-2 tokenizer. In our experiments, we train with $10000$ URF $4$-addition samples.
We choose GPT-2 with key-only position embeddings as our model. For the RPE and RPE-Square settings, we augment the GPT-2 model with RPE and RPE-Square, respectively. The implementation is adapted from HuggingFace [23].
We train the models by AdamW, with the initial learning rate $0.0005$, the weight decay $1.0$, and the cosine scheduler. The warmup ratio is $0.05$. The gradient accumulation steps is $2$. The per-device training batch size is $128$. We train the models for $20000$ steps and $200000$ steps. The experiments are run on a server with Ubuntu. All the models are trained on two NVIDIA GeForce RTX 3090 GPUs.